# Image Classification in High-Energy Physics: A Comprehensive Survey of Applications to Jet Analysis

### v2

@@ Abstract
Nowadays, there has been a growing trend in the fields of high-energy physics (HEP) in its both parts experimental and phenomenological studies, to incorporate machine learning (ML) and its specialized branch, deep learning (DL). This review paper provides a thorough illustration of these applications using different DL approaches. The first part of the paper examines the basics of various particle physics types and sets up guidelines for assessing particle physics alongside the available learning models. Next, a detailed classification is provided for representing the jet images that are reconstructed in high energy collisions mainly with proton-proton collisions at well defined beam energies, covering various datasets, preprocessing techniques, and feature extraction and selection methods. The presented techniques can be applied to future hadron-hadron colliders (HLC) such as high luminosity LHC (HL-HLC) and future circular collider - hadron-hadron (FCC-hh). Next, the authors explore a number of AI models analysis designed specifically for images in HEP. We additionally undertake a closer look at the classification associated with images in hadron collisions, with an emphasis on Jets. In this review, we look into various state-of-the-art (SOTA) techniques in ML and DL, examining their implications for HEP demands. More precisely, this discussion tackles various applications in extensive detail, such as Jet tagging, Jet tracking, particle classification, and more. The review concludes with an analysis of the current state of HEP, using DL methodologies. It covers the challenges and potential areas for future research that will be illustrated for each application.
@@ 1 Introduction
High-energy physics (HEP) is an attracting and delicate branch of physics that manifests at the microscopic scale and which explores the fundamental building blocks of the universe and forces that govern their interactions at incredibly high energies under extremely intense conditions [1, 2]. In this field many sophisticated instruments and tools with large particle accelerators, like the current CERN-LHC (located near the French and Swiss border), to study matter at energy levels that are otherwise unattainable to reach with conventional methods. These gigantic machines accelerate subatomic particles at nearly the speed of light and then smash them together, creating energy densities analogous of the early moments after the Big Bang [3, 4]. By studying the collisions generated in these accelerators setups, it could be possible to track and evaluate rare particles that have a very short life time. This important study with the accumulated big data at higher collider luminosity values offers an improved understanding of the basic anatomy of different physics process and their topologies [5, 6].
@@
The standard model (SM) is the present theoretical framework that describes the elementary particles and their interactions [7]. Despite its tremendous success in explaining many phenomena in nature, several mysteries remain unsolved, such as matter antimatter asymmetry, the nature of Dark Matter(DM), the neutrino mass and the hierarchy problem and many other open questions [8]. Furthermore, it is worth noting that besides its deep investigation about the Universe puzzles, HEP has demonstrated significant practical utility when used with advanced technologies [9]. As a matter of fact, the development of many techniques and technologies in this sector has driven notable progress in medical imaging [10, 11], radiation therapy [12, 13], and materials research [14, 15].
@@
The data acquisition system of large hadron collider (LHC) stores the data on tape using grid computing facilities, it can be disseminated for offline analysis aimed at extracting information concerning particle trajectories formed within the detectors. These trajectories contain concealed details about numerous particle characteristics. Computer vision techniques become relevant and play a crucial role during the analysis of offline data. Specifically, in the realm of HEP data analysis, machine learning (ML) algorithms have found success, leading to significant enhancements in event classification performance when contrasted with traditional methods rooted in expert understanding. Techniques like boosted decision trees (BDT), shallow neural networks, and similar approaches have been employed in HEP data analysis. More recently, deep neural network (DNN) or deep learning (DL) have gained widespread adoption due to their applicability to intricate data structures such as images, videos, natural language, or sensor data. There are ongoing investigations into applying DNNs for analyzing granular details like particle positions and momentum as they traverse the detector. This has shown increased effectiveness in selecting signal events compared to ML algorithms employing conventional feature variables rooted in physics knowledge [16].
@@ 1.1 Motivation
In the light of the aforementioned phenomena and challenges, treating tracks / Jets in HEP as image-like data for processing and analysis is a useful approach, especially when dealing with the output from particle detectors. Hence, ML and DL play vital roles in HEP experiments. They serve the following purposes: i) Identifying and classifying particles by analyzing their tracks and energy deposits in detectors, thereby enhancing precision and identification speed, ii) assisting in the accurate reconstruction of particle tracks from detector data, particularly in complex environments with numerous particles and interactions, iii) enabling efficient data analysis schemes, one can sift through extensive datasets to pinpoint rare or noteworthy events or particles, iv) detecting anomalies or unexpected patterns in the recorded data, which could potentially signify the existence of new particles and physics beyond the SM, among other applications. These contributions underscore the significance of ML and DL in advancing HEP research topics.
@@ 1.2 Related work
In recent years, there has been a surge in reviews addressing various aspects of HEP [17, 18, 19, 20, 21]. The review presented in [17] delved into the realm of supervised deep learning applied to high-energy phenomenology, discussing specific use cases such as employing ML to explore new physics parameter spaces and utilizing graph neural networks for particle production and energy measurements at the LHC. Meanwhile, paper [18] provided an overview of the initial forays into quantum ML in the context of HEP and offered insights into potential future applications. In [19], an array of novel tools relevant to HEP were introduced, complete with assessments of their performance, though there was limited discussion about future prospects. Lastly, the reviews [20, 21] comprehensively examined both theoretical and experimental aspects of Jets such as triggering, data acquisition systems, propagation, interactions, and related phenomena in HEP.
Table 1 assesses how the proposed review aligns with previous research in the field of HEP. Based on the assessment, it appears that our proposed review aims to comprehensively cover a wide range of topics related to the ML and DL-based in HEP, including Jet preliminaries, taxonomy of HEP, available Jet datasets, Jet image preprocessing, quantum ML, DL models for Jet images, classification techniques, Jet image DL applications, and research gaps/future directions. This suggests that the proposed review aims to provide a comprehensive overview of the current state of research in HEP and potential avenues for future work.
@@
Reference Paper type Publication year Jet prelimanaries Taxonomy of HEP Available Jet datasets Jet images pre-process Quantium ML for Jet images DL models for Jet images Jet image DL classif. techniques Jet image DL apps Research gaps and future direction
@@ 1.3 Contribution and survey structure
The objective of this survey is to provide a robust foundation for both HEP researchers aiming to grasp the principles of deep learning and its applications within the HEP domain, and for computer science researchers familiar with artificial intelligence (AI) seeking insights into the fundamental features and prerequisites essential for constructing a robust AI model tailored specifically for HEP, employing Jet images. To achieve this goal, our contribution is encapsulated in the following key points:
@@
  • The taxonomy of ML and DL-based techniques in HEP for analyzing Jet images, along with their respective preprocessing and feature extraction methodologies, is thoroughly explored.
  • The widely adopted AI models designed for analyzing HEP Jet images, along with their descriptive layered architectures, are extensively elaborated upon. Furthermore, their performance metrics are summarized and compared.
  • Different state-of-the art (SOTA) methods are clustered based on the AI techniques employed and comprehensively reviewed accordingly. Additionally, the exploration of AI-based applications in HEP Jet images is thoroughly detailed.
  • Future directions and outlooks are explored, which aims to offer researchers insights into existing research gaps and areas within AI concepts and fields that remain unexplored in AI-based Jet images.
The structure of this paper is as follows: Section 2 presents the preliminaries necessary for understanding Jet images. In Section 3, the representation of Jet images in HEP is discussed. Section 4 provides a summary of the available AI models for analyzing HEP Jet images. Section 5 reviews AI-based methods for Jet images and categorizes them based on whether they employ ML or DL techniques. Section 6 showcases various AI-based applications of Jet images. Section 7 highlights the gaps and areas that remain unexplored in AI-based Jet image analysis, encompassing both techniques and applications. Finally, Section 8 concludes the survey.
@@ 2.1 Types of particles
W and Z bosons: are important closely related particles described by the SM of particle physics. They are together known as the weak bosons or more generally as the intermediate vector bosons and plays a significant role in the weak nuclear force, which is responsible for certain types of specific interactions and radioactive decay. The existence and properties of the $Z$ boson, along with the $W$ bosons, provided strong support for the electroweak theory and the SM as a whole. However, as with the $W$ boson, the Standard Model has limitations and does not explain all aspects of particle physics, such as gravity, dark matter, and the hierarchy of particle masses. Here are some key points about the $W$ and $Z$ bosons:
@@
The top quark: is one of the heavy fundamental particles described by the Standard Model. It holds a special place in particle physics due to its extremely large mass and its role in various processes involving high-energy collisions. Here are some key points about the top quark [24]:
@@
  • Weak Decays: Due to its high mass, the top quark is relatively short-lived and decays before it can form bound states with other quarks to create hadrons. It decays primarily through weak interaction, one of the fundamental forces described by the Standard Model.
@@ 2.2 Performance measures
Table 2: An overview of the metrics employed to evaluate performance in particle physics.
Metric Formula Description
Classic metrics Energy loss ratio $\chi_{jh}\equiv\frac{E_{f}^{h}}{E_{i}^{h}}$ Determine, on a Jet-by-Jet basis, the amount of energy loss.The notation "jh" denotes the energy of the Jet at the hadronic scale. Here, $E_{f}^{h}$ represents the $p_{T}$ of a specific Jet when interacting with a medium, whereas $E_{i}^{h}$ indicates the $p_{T}$ of the identical Jet in the absence of any medium.
Traversed path-length $L=\frac{\sum_{i\in Jet}p_{T,i}L_{i}}{\sum_{i\in Jet}p_{T,i}}$ Offers significant insights closely related to the alterations and energy dissipation encountered by a Jet. This involves calculating the path length, $L$, traversed by a parton Jet, which is derived from the sum of the lengths of the Jet’s constituents at the partonic level $L_{i}$, weighted by their $p_{T}$.
Jet axis distance $\Delta R=\sqrt{(\Delta\eta)^{2}+(\Delta\phi)^{2}}$ The constituents of a Jet are predominantly confined within a conical shape, characterized by a separation from the Jet axis of $\Delta R=0.5$. This separation $\Delta R$ encompasses both $\Delta\eta$, the variation in pseudo-rapidity, and $\Delta\phi$, the deviation in azimuthal angle relative to the Jet axis.
Efficiency and
Miss-efficiency rates
Eff = $\frac{N_{top}^{top}}{N_{top}},\quad\text{Mistag}=\frac{N_{QCD}^{top}}{N_{QCD}}$ Defining efficiency and mis-tag rates is crucial for assessing the performance of the artificial neural network (ANN) tagger. Where $N_{\text{top}}$ and $N_{\text{QCD}}$ represent the total jet count in the top and QCD jet samples, respectively, and $N_{top}^{QCD}$ signifies the number of jets in sample $top$ tagged as type $QCD$.
FPR and TPR $\mathrm{\frac{FP}{FP+TN}},\mathrm{\frac{TP}{TP+FN}}$ The FPR, is the ratio (or percentage) of the background signal that are incorrectly identified as containing Jet. The TPR. is the ratio (or percentage) of the Jet signal that is correctly identified as Jet (particle).
AUC $\int_{0}^{1}\mathrm{TPR}\>d(\mathrm{FPR})$ The area beneath the ROC curve is represented. It delivers a singular numeric score reflecting the cumulative effectiveness of the classification technique. An elevated AUC score signifies superior performance, with the ideal score being 1.
@@
Deep learning metrics Accuracy $\mathrm{\frac{TP+TN}{TN+FN+TP+FP}}$ The accuracy is the ratio (or percentage) of correctly detected instances of Jet in the signal. A high accuracy indicates that the classification algorithm is more effective in detecting Jet than background.
MSE $E(\theta)=\frac{1}{N}\sum_{i=1}^{N}(P_{b}^{i}(\theta)-T^{i})^{2}$ The training procedure seeks to discover the model parameter values denoted as $\theta$, which minimize the loss function known as MSE. Where $N$ is the number of training Jets, $P_{b}^{i}$ and $T^{i}$ is the predicted and target probabilities, respectively, for the $i$-th Jet.
@@
F-measure $2.\mathrm{\frac{Precision\times Recall}{Precision+Recall}}$ Represents the harmonic mean between precision and recall metrics. This measure is applied to assess the comprehensive efficacy of the classification algorithm in identifying or tagging Jets.
Joint and conditional distributions $P(v\backslash O)\neq P(v)\hskip 8.5359pt\forall O$ If the probability distribution of the physics variable $v$, given the neural network output $O$, is influenced by the neural network output itself, then we interpret this as an indication that the network has acquired knowledge about this physics characteristic.
FST $\frac{1}{{(1+\sum_{0}^{n}|z_{n}|^{2})^{2}}}\sum_{i=0}^{n}|dz_{i}|^{2}$ The FST is an invariant metric tensor that can be used to describe distances between quantum states [27, 28]. $z_{i}$ are complex coordinates, and $|dz_{i}|$ represents the absolute value of the differential of $z_{i}$.
AMS $\sqrt{2\left((S+B+B_{r})\ln\left(1+\frac{S}{B+B_{r}}\right)-S\right)}$ The effectiveness of a method in categorizing test data into signal or background groups is assessed by employing approximate median significance (AMS) metric [19]. $S$ is the sum of weights of true positives; $B$ is the sum of weights of false positives; $B_{r}$ is the regularization constant.
3 HEP Jet image representation

### v3

@@ Abstract
Nowadays, there has been a growing trend in the field of high-energy physics (HEP), in both its experimental and phenomenological studies, to incorporate machine learning (ML) and its specialized branch, deep learning (DL). This review paper provides a thorough illustration of these applications using different ML and DL approaches. The first part of the paper examines the basics of various particle physics types and establishes guidelines for assessing particle physics alongside the available learning models. Next, a detailed classification is provided for representing Jets that are reconstructed in high-energy collisions, mainly in proton-proton collisions at well-defined beam energies. This section covers various datasets, preprocessing techniques, and feature extraction and selection methods. The presented techniques can be applied to future hadron-hadron colliders (HHC), such as the high-luminosity LHC (HL-LHC) and the future circular collider - hadron-hadron (FCC-hh). The authors then explore several AI techniques analyses designed specifically for both image and point-cloud (PC) data in HEP. Additionally, a closer look is taken at the classification associated with Jet tagging in hadron collisions. In this review, various state-of-the-art (SOTA) techniques in ML and DL are examined, with a focus on their implications for HEP demands. More precisely, this discussion addresses various applications in extensive detail, such as Jet tagging, Jet tracking, particle classification, and more. The review concludes with an analysis of the current state of HEP using DL methodologies. It highlights the challenges and potential areas for future research, which are illustrated for each application.
@@
Abbreviations list
AE
auto-encoder
AI
artificial intelligence
ANN
artificial neural network
AUC
area under curve
BDT
boosted decision trees
BIP
boost-invariant polynomial
BRNN
bidirectional RNN
CERN
conseil Européen pour la recherche nucléaire
CG
Clebsch-Gordan
CGENN
Clifford group equivariant neural networks
CMS
compact muon solenoid
CMSSM
constrained minimal super-symmetric standard model
CNN
convolutional neural network
DL
deep learning
DLA
dataset link availability
DNN
deep neural network
EG
event generator
EGNN
equivariant graph neural networks
FAR
false alarm rate
FGSM
fast gradient sign method
FL
federated learning
FPR
false-positive rate
FST
Fubini-study tensor
GAN
generative adversarial network
GMM
Gaussian mixture model
GNN
graph neural network
GPU
graphics processing unit
GRU
gated recurrent unit
HCAL
hadronic calorimete
HEP
high-energy physics
HPC
high-performance computing
KNN
K-nearest neighbors
L-GATr
Lorentz geometric algebra Transformer
LGEB
Lorentz group equivariant block
LGN
Lorentz group network
LHC
large hadron collider
LLM
large language model
LSTM
long short-term memory
MIParT
more-interaction particle Transformer
ML
machine learning
MLP
multi-Layer perceptron
MSE
mean squared error
PC
point cloud
PCA
principal component analysis
PCT
point cloud Transformer
$p_{T}$
transverse momentum
PFN
particle flow network
PFN-ID
particle flow network with ID
PLP
primary lund plane
P-MHA
pairwise multi-head attention
QCD
quantum chromodynamics
QGP
quark-gluon plasma
QML
quantum machine learning
RecNN
recursive neural network
ReLU
rectified linear unit
RL
reinforcement learning
RNN
recurrent neural network
ROC
receiver operating characteristic
SHAP
shapley additive explanations
SM
standard model
SOTA
state-of-the art
SVM
support vector machine
TL
transfer learning
TPR
true positive rate
t-SNE
t-distributed stochastic neighbor embedding
UMAP
uniform manifold approximation and projection
ViT
vision Transformers
@@ 1 Introduction
HEP is an attracting and delicate branch of physics that manifests at the microscopic scale and which explores the fundamental building blocks of the universe and forces that govern their interactions at incredibly high energies under extremely intense conditions [1, 2]. In this field many sophisticated instruments and tools with large particle accelerators, like the current CERN-LHC (located near the French and Swiss border), to study matter at energy levels that are otherwise unattainable to reach with conventional methods. These gigantic machines accelerate subatomic particles at nearly the speed of light and then smash them together, creating energy densities analogous of the early moments after the Big Bang [3, 4]. By studying the collisions generated in these accelerators setups, it could be possible to track and evaluate rare particles that have a very short life time. This important study with the accumulated big data at higher collider luminosity values offers an improved understanding of the basic anatomy of different physics process and their topologies [5, 6].
@@
The standard model (SM) is the present theoretical framework that describes the elementary particles and their interactions [7]. Despite its tremendous success in explaining many phenomena in nature, several mysteries remain unsolved, such as matter antimatter asymmetry, the nature of dark matter (DM), the neutrino mass and the hierarchy problem and many other open questions [8]. Furthermore, it is worth noting that besides its deep investigation about the Universe puzzles, high-energy physics (HEP) has demonstrated significant practical utility when used with advanced technologies [9]. As a matter of fact, the development of many techniques and technologies in this sector has driven notable progress in medical imaging [10, 11], radiation therapy [12, 13], and materials research [14, 15].
The data acquisition system of large hadron collider (LHC) stores the data on tape using grid computing facilities, it can be disseminated for offline analysis aimed at extracting information concerning particle trajectories formed within the detectors. These trajectories contain concealed details about numerous particle characteristics. Jets are reconstructed by combining information from multiple detector subsystems, primarily calorimeters and trackers. The calorimeters (electromagnetic and hadronic) play a central role by capturing the energy deposits from both neutral and charged particles. These deposits are clustered using algorithms such as anti-$k_{t}$, which group the energy into Jets based on angular proximity in $(\eta,\phi)$-space. While tracking systems provide detailed momentum and charge information for individual charged particles, they cannot detect neutral particles, such as photons or neutrons. Therefore, the calorimeter serves as the primary tool for measuring the total energy of the Jet. This reconstruction process ensures that Jets are defined as comprehensive objects representing the full range of particle constituents, crucial for subsequent analyses in HEP experiments [16].
@@
Computer vision techniques become relevant and play a crucial role during the analysis of offline data. Specifically, in the realm of HEP data analysis, machine learning (ML) algorithms have found success, leading to significant enhancements in event classification performance when contrasted with traditional methods rooted in expert understanding. Techniques like boosted decision trees (BDT), shallow neural networks, and similar approaches have been employed in HEP data analysis. More recently, deep neural network (DNN) or deep learning (DL) have gained widespread adoption due to their applicability to intricate data structures such as images, videos, natural language, or sensor data. There are ongoing investigations into applying DNNs for analyzing granular details like particle positions and momentum as they traverse the detector. This has shown increased effectiveness in selecting signal events compared to ML algorithms employing conventional feature variables rooted in physics knowledge [17].
@@ 1.1 Motivation
In the light of the aforementioned phenomena and challenges, treating tracks / Jets in HEP as image or point cloud (PC)-like data for processing and analysis is a useful approach, especially when dealing with the output from particle detectors. Hence, ML and DL play vital roles in HEP experiments. They serve the following purposes: i) Identifying and classifying particles by analyzing their tracks and energy deposits in detectors, thereby enhancing precision and identification speed, ii) assisting in the accurate reconstruction of particle tracks from detector data, particularly in complex environments with numerous particles and interactions, iii) enabling efficient data analysis schemes, one can sift through extensive datasets to pinpoint rare or noteworthy events or particles, iv) detecting anomalies or unexpected patterns in the recorded data, which could potentially signify the existence of new particles and physics beyond the SM, among other applications. These contributions underscore the significance of ML and DL in advancing HEP research topics.
@@ 1.2 Related work
In recent years, there has been a surge in reviews addressing various aspects of HEP [18, 19, 20, 21]. The review presented in [18] delved into the realm of supervised DL applied to high-energy phenomenology, discussing specific use cases such as employing ML to explore new physics parameter spaces and utilizing graph neural networks for particle production and energy measurements at the LHC. Meanwhile, paper [19] provided an overview of the initial forays into quantum ML in the context of HEP and offered insights into potential future applications. In [20], an array of novel tools relevant to HEP were introduced, complete with assessments of their performance, though there was limited discussion about future prospects. Lastly, the review [21] comprehensively examined both theoretical and experimental aspects of Jets such as triggering, data acquisition systems, propagation, interactions, and related phenomena in HEP.
Table 1 assesses how the proposed review aligns with previous research in the field of HEP. Based on the assessment, it appears that our proposed review aims to comprehensively cover a wide range of topics related to the ML and DL-based in HEP, including Jet preliminaries, taxonomy of HEP, available Jet datasets, Jet tagging preprocessing, quantum ML, DL models for Jet tagging, classification techniques, Jet tagging DL applications, and research gaps/future directions. This suggests that the proposed review aims to provide a comprehensive overview of the current state of research in HEP and potential avenues for future work.
@@
Reference Paper type Publication year Jet prelimanaries Taxonomy of HEP Jet Available Jet datasets
and tools
Jet tagging pre-process Quantium ML for HEP Jet classification ML and DL models for Jet classification Transformers for Jet classification ML and DL-based Jet classif. techniques AI-based Jet apps Research gaps and future direction
@@ 1.3 Contribution and survey structure
The objective of this survey is to provide a robust foundation for both HEP researchers aiming to grasp the principles of DL and its applications within the HEP domain, and for computer science researchers familiar with artificial intelligence (AI) seeking insights into the fundamental features and prerequisites essential for constructing a robust AI model tailored specifically for HEP, employing Jet images and PC. To achieve this goal, our contribution is encapsulated in the following key points:
@@
  • The taxonomy of ML and DL-based techniques in HEP for analyzing Jet images and PC, along with their respective preprocessing and feature extraction methodologies, is thoroughly explored.
  • The widely adopted AI models designed for analyzing HEP Jet tagging, along with their descriptive layered architectures, are extensively elaborated upon. Furthermore, their performance metrics are summarized and compared.
  • Different state-of-the art (SOTA) methods are clustered based on the AI techniques employed and comprehensively reviewed accordingly. Additionally, the exploration of AI-based applications in HEP Jet classification is thoroughly detailed.
  • Future directions and outlooks are explored, which aims to offer researchers insights into existing research gaps and areas within AI concepts and fields that remain unexplored in AI-based Jet images and PC.
The structure of this paper is as follows: Section 2 presents the preliminaries necessary for understanding Jet images and PC. In Section 3, the representation of Jet in DL-based HEP is discussed. Section 4 provides a summary of the most available ML or DL models for analyzing HEP Jet tagging. Section 5 showcases various AI-based applications of Jet tagging. Section 6 highlights the gaps and areas that remain unexplored in AI-based Jet analysis, encompassing both techniques and applications. Finally, Section 7 concludes the survey.
@@ 2.1 Types of particles
W and Z bosons: are important closely related particles described by the SM of particle physics. They are together known as the weak bosons or more generally as the intermediate vector bosons and plays a significant role in the weak nuclear force, which is responsible for certain types of specific interactions and radioactive decay. The existence and properties of the $Z$ boson, along with the $W$ bosons, provided strong support for the electroweak theory and the SM as a whole. However, as with the $W$ boson, the SM has limitations and does not explain all aspects of particle physics, such as gravity, dark matter, and the hierarchy of particle masses. Here are some key points about the $W$ and $Z$ bosons:
@@
The top quark: is one of the heavy fundamental particles described by the SM. It holds a special place in particle physics due to its extremely large mass and its role in various processes involving high-energy collisions. Here are some key points about the top quark [25]:
@@
  • Weak decays: Due to its high mass, the top quark is relatively short-lived and decays before it can form bound states with other quarks to create hadrons. It decays primarily through weak interaction, one of the fundamental forces described by the SM.
@@
2.2 Key concepts of ML-based HEP
When discussing ML and its subset DL in HEP, maintaining uniform and precise terminology is crucial for clear communication. Supervised learning, for instance, refers to training models using labeled datasets, where the model learns to map input features to known outputs, such as identifying particles or classifying Jets based on their physical properties. In contrast, unsupervised learning involves identifying patterns or structures in data without predefined labels, often used in anomaly detection or clustering in particle physics. Feature selection is an essential process that focuses on choosing the most informative input features—such as track momentum, calorimeter energy deposits, and hit patterns in detectors—thereby improving the performance and efficiency of ML models by reducing dimensionality and computational load.
The growing adoption of DL techniques, like convolutional neural networks and graph neural networks, has revolutionized analyses in HEP. These methods rely on different types of layers and architectures designed to handle the complexity and scale of particle physics data. Convolutional layers in CNNs, for instance, are particularly effective at detecting patterns in images or PCs, by learning local features. These layers operate by applying convolutional filters to input data, extracting hierarchical patterns, which are then pooled to reduce dimensionality. Pooling layers, such as max-pooling, downsample the spatial dimensions of the data, retaining the most important features while reducing computational cost. This structure allows CNNs to efficiently process large-scale data and is widely used in Jet classification and particle identification tasks. Further advancements include the use of EdgeConv layers in GNNs [28], where the network learns the relationships between particles represented as nodes in a graph. In these models, the EdgeConv block aggregates local particle information, capturing spatial relationships and interactions based on particle kinematics and connectivity, which are essential for Jet tagging. The use of global average pooling in these models helps aggregate information from individual particles, producing a global representation of the Jet that can then be used for classification or regression tasks. Dense layers (also known as fully connected layers) play a critical role in transforming high-level features learned by convolutional and graph-based layers into a final prediction. Dense layers are used in DNNs, CNNs, GNNs, and others, after the feature extraction phase, where the output of the convolutional or graph layers is flattened into a one-dimensional vector and passed through one or more fully connected layers. These layers allow the network to combine the learned features in a non-linear way, making complex decisions such as event classification, particle identification, or regression for Jet properties. The dense layer’s ability to connect all input neurons to all output neurons allows the model to capture intricate relationships between features, making it highly effective for tasks like anomaly detection, signal classification, and event reconstruction in HEP.
An essential innovation in modern DL is the Attention layer [29], a core layer in building Transformers, enables the model to focus on the most relevant parts of the input data. Attention mechanisms are particularly useful in scenarios where certain elements in a sequence (or graph) are more important for the task than others. In particle physics, this could involve focusing on particular particle interactions or energy deposits in Jets. The scaled dot-product Attention mechanism, used in Transformer models, computes attention scores for each pair of input elements [30]. The attention output $\text{Attention}(Q,K,V)$ is calculated as follows:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$ (1)
Where $Q$, $K$, and $V$ represent the query, key, and value matrices, respectively, and $d_{k}$ is the dimension of the key vectors. The softmax function normalizes the attention scores, allowing the model to weigh the importance of different elements in the input sequence. This mechanism enables the model to prioritize relevant information, improving the accuracy of particle event classification, Jet tagging, and anomaly detection, particularly when the input data has complex dependencies or long-range interactions between particles.
@@ 2.3 Performance measures
Table 2: An overview of the metrics employed to evaluate performance in ML and DL-based HEP.
Metric Formula C/R Description
[…]
@@
Accuracy $\mathrm{\frac{TP+TN}{TN+FN+TP+FP}}$ C The accuracy is the ratio (or percentage) of correctly detected instances of Jet in the signal. A high accuracy indicates that the classification algorithm is more effective in detecting Jet than background.
[…]
@@
F1-score $2.\mathrm{\frac{Precision\times Recall}{Precision+Recall}}$ C Represents the harmonic mean between precision and recall metrics. This measure is applied to assess the comprehensive efficacy of the classification algorithm in identifying or tagging Jets.
Abbreviations: Classification or regression (C/R);
3 HEP Jet representation